OpenStudy (anonymous):

Convert the equation r=-6 csc theta to rectangular form.

zepdrix (zepdrix):

\[\Large\bf\sf r=-6 \csc \theta\]\[\Large\bf\sf r=-6\frac{1}{\sin \theta}\]Multiply both sides by sin(theta), then use your rectangular/polar identities,\[\Large\bf\sf r \sin \theta =x,\qquad\qquad r \cos \theta=y\]

zepdrix (zepdrix):

Woops I wrote those backwards,\[\Large\bf\sf r \sin \theta =y,\qquad\qquad r \cos \theta = x\]